OpenStudy (anonymous):

How do you find the slope and y-intercept of 3x-8y=-24

OpenStudy (katrinakaif):

You try: Turn it into a y=mx + b format

OpenStudy (anonymous):

Well i got -8y=3x+(-24) but i got stuck here

OpenStudy (katrinakaif):

Okay, now divide it by -8

OpenStudy (anonymous):

Divide 3x+(-24)?

OpenStudy (katrinakaif):

Wait..

OpenStudy (katrinakaif):

when you bring 3x to the other side, you are substracting it.. therefore your final equation should be -8y = -3x - 24 correct?

OpenStudy (katrinakaif):

Now divide that by -8

OpenStudy (anonymous):

Divede -3x-24 by -8?

OpenStudy (katrinakaif):

Yup so -3/-8 x and -24/-8

OpenStudy (katrinakaif):

Lol yeah :P -3/-8 becomes 3/8 and -24/-8 becomes + 3

OpenStudy (anonymous):

Alright, so what do i do from there?

OpenStudy (katrinakaif):

Okay, so your equation is y = 3/8x + 3 and is in y= mx + b format Now m = slope so what is your slope?

OpenStudy (anonymous):

0.375?

OpenStudy (katrinakaif):

or simply 3/8

OpenStudy (anonymous):

Oh alright haha and would the y-intercept be 3?

OpenStudy (katrinakaif):

Yes! Great work :D

OpenStudy (anonymous):

Thanks :D

OpenStudy (katrinakaif):

Anytime =)